OpenStudy (anonymous):

The wrestling team runs 10 laps at the beginning of each practice. for every minute a team member is late for practice, he must run additional laps. Chen was 5 minutes late for practice and had to run a total of 25 laps. write a linear equation that gives the number of laps run, Y, in terms of the number of minutes late for practice, x.

OpenStudy (anonymous):

can anybody answer this question

OpenStudy (anonymous):

10 is the set number so if a person was 0 min late we know that he runs 10 laps. he was 5 min late and had to run a total of 25 laps so 10 (inital) + 5min(x laps/min) = 25 10 +5x =25 x=3 laps per 1 min late so y =laps to run y=10 + 3x